# SWIFT: Semantic Watermarking for Image Forgery Thwarting

### v1

@@ Abstract
Index Terms:
Watermarking, Image authentication, Semantic information
@@ I Introduction
  • $-\mathbb{R}$ : Inspired by HiDDeN [5], we propose an encoder-decoder network architecture jointly trained to embed and extract high-dimensional unit-norm vectors in images.
@@ III Method
Figure 1: Overview of SWIFT. To ensure the integrity of an image, Alice first leverages the message layer to tailor a semantic representation of the image, in our case a caption compressed in a lossless fashion. The resulting bit stream is fed to the TCCSK modulation layer thus enabling security and confidence (see Sect. IV-C), and then to the watermarking layer based on our $-\mathbb{R}$ encoder-decoder neural network. At reception, Bob executes the inverse process with a secret key and obtains both the caption and the p-value $\rho$. If $\rho$ is low enough, Bob can entrust the decoded caption and compare the image he received with the caption, enabling comparison between a proxy of the original image and the received one.
@@ III-B The watermarking layer
Figure 3: $-\mathbb{R}$ architecture. We use a L2 norm to control the watermark power to enforce a target PSNR on both the watermark signal in the embedder and on the decoded vector in the extractor.
@@
Fig. 3 depicts our $-\mathbb{R}$ architecture, resulting from several modifications of HiDDeN. The input data works with unit vectors in high-dimensional real space instead of binary messages. Specifically, we draw random samples $X$ uniformly distributed on the surface of the unit hypersphere in $\mathbb{R}^{D}$ with $D=256$ :
@@ III-C2 Modulating variable-rate signals
[…]
@@ IV-C Confidence metric
After the TCCSK modulation, a message $M$ is encoded into a vector $X$ on the 256-d hypersphere. Given an encoded message $X$, due to transforms during transmission, $Y$ is the noisy extracted vector. At inference, $Y$ is decoded by the TCCSK demodulation into $\hat{M}$. $X$ is unknown, but let $\hat{X}$ be the perfect encoding of $\hat{M}$ and $C$ the cosine similarity computed between $\hat{X}$ and $Y$ (see Fig. 5). We define our pratical confidence metric as $\rho$ (7). We assume that $Y$ close to the perfect representation $\hat{X}$ of the decoded message $\hat{M}$ means that $Y$ is also close to the (unknown) originally embedded vector $X$. Thus, a low $\rho$ would entail a low number of errors at the decoding step, as confirmed by the Pearson correlation coefficient of $-0.89$ computed between $8,000$ $\rho$ values and corresponding MRR. Tab. II shows the MRR on watermarked images under several scenarios: watermarked images, benign attacks and malign attacks, and three $\rho$-thresholds. Only messages with $\rho<\rho$-threshold should be trusted. An adequate threshold ensures the perfect extraction, meaning $100\%$ MRR, of the watermarked caption. Note that $\rho$ also refers to the error probability on non-watermarked images: the higher it is, the more a non-watermarked image could be flagged as watermarked.
A confidence metric could also be used in a multi-bit setting with fixed-length code of 64 bits. Padding a $n$-bit message with $64-n$ equiprobable random bits drawn from a synchronous source shared by Alice and Bob (e.g. via a secret key $K$) allows to carry confidence information, at the expense of capacity. Indeed, Bob can check the padding bits and discard any message not matching the expected sequence. In this case, assuming the multi-bit system outputs random codes uniformly under $\mathcal{H}_{0}$, there is still a $2^{-64+n}$ probability of ending up with the expected sequence by chance, giving $\rho\geq 2^{-64}\simeq 5e-20$. Our method benefits from comparable confidence, with $\rho=1.4e-16$ to be protected against all attacks, along with a much greater capacity.
@@ V Conclusion
In this work, we present a novel way to assert the integrity of an image by the relevant use of watermarking as a covert communication channel. Moreover we provide a definition of the image semantics, through its caption, to the recipient of the message. By using an LLM combined with an arithmetic encoder to compress the caption, the limited capacity of $-\mathbb{R}$ to convey information is optimized.

### v2

@@ I Introduction
  • Hide-$\mathbb{R}$ : Inspired by HiDDeN [5], we propose an encoder-decoder network architecture jointly trained to embed and extract high-dimensional unit-norm vectors in images.
@@ III Method
Figure 1: Overview of SWIFT. To ensure the integrity of an image, Alice first leverages the message layer to tailor a semantic representation of the image, in our case a caption compressed in a lossless fashion. The resulting bit stream is fed to the TCCSK modulation layer thus enabling security and confidence (see Sect. IV-C), and then to the watermarking layer based on our Hide-$\mathbb{R}$ encoder-decoder neural network. At reception, Bob executes the inverse process with a secret key and obtains both the caption and the p-value $\rho$. If $\rho$ is low enough, Bob can entrust the decoded caption and compare the image he received with the caption, enabling comparison between a proxy of the original image and the received one.
@@ III-B The watermarking layer
Figure 3: Hide-$\mathbb{R}$ architecture. We use a L2 norm to control the watermark power to enforce a target PSNR on both the watermark signal in the embedder and on the decoded vector in the extractor.
@@
Fig. 3 depicts our Hide-$\mathbb{R}$ architecture, resulting from several modifications of HiDDeN. The input data works with unit vectors in high-dimensional real space instead of binary messages. Specifically, we draw random samples $X$ uniformly distributed on the surface of the unit hypersphere in $\mathbb{R}^{D}$ with $D=256$ :
@@ III-C2 Modulating variable-rate signals
In the previous section, Bob needs to test $2^{N}$ keys which becomes intractable as $N$ grows. An alternative is to generate multiple carriers from the same pseudo-random generator seeded by secret key $K$, and to combine them to produce $X\in\mathbb{R}^{D}$. This paper uses truncated cyclic shift keying (TCCSK) modulation [25] for its improved performance in the additive white Gaussian noise (AWGN) channel compared to Binary Phase Shift Keying (BPSK) (see Fig. 4). The message $M$ is split into $T$ equal blocks of length $L$, padding the last block with 0 if necessary. Then the $j$-th block is encoded from a carrier $Z_{j}\in\mathbb{R}^{2^{L}}\sim\mathcal{N}(0,I_{2^{L}}),j\in\llbracket 0\mkern 1.5mu..\mkern 1.5muT-1\rrbracket$ by cyclically shifting it by a number equal to the value represented by the $L$-bit sub-message $M_{j}\coloneqq M_{\llbracket jL\mkern 1.5mu..\mkern 1.5mu(j+1)L\rrbracket}$. The carriers are truncated to dimension $D$, normalized, summed, and normalized again for transmission:
@@ IV-C Confidence metric
After the TCCSK modulation, a message $M$ is encoded into a vector $X$ on the 256-d hypersphere. Given an encoded message $X$, due to transforms during transmission, $Y$ is the noisy extracted vector. At inference, $Y$ is decoded by the TCCSK demodulation into $\hat{M}$. $X$ is unknown, but let $\hat{X}$ be the perfect encoding of $\hat{M}$ and $C$ the cosine similarity computed between $\hat{X}$ and $Y$ (see Fig. 5). We define our pratical confidence metric as $\rho$ (7). We assume that $Y$ close to the perfect representation $\hat{X}$ of the decoded message $\hat{M}$ means that $Y$ is also close to the (unknown) originally embedded vector $X$. Thus, a low $\rho$ would entail a low number of errors at the decoding step, as confirmed by the Pearson correlation coefficient of 0.89-0.89 computed between $8,000$ $\rho$ values and corresponding MRR. Tab. II shows the MRR on watermarked images under several scenarios: watermarked images, benign attacks and malign attacks, and three $\rho$-thresholds. Only messages with $\rho<\rho$-threshold should be trusted. An adequate threshold ensures the perfect extraction, meaning $100\%$ MRR, of the watermarked caption. Note that $\rho$ also refers to the error probability on non-watermarked images: the higher it is, the more a non-watermarked image could be flagged as watermarked.
A confidence metric could also be used in a multi-bit setting with fixed-length code of 64 bits. Padding a $n$-bit message with 64n64-n equiprobable random bits drawn from a synchronous source shared by Alice and Bob (e.g. via a secret key $K$) allows to carry confidence information, at the expense of capacity. Indeed, Bob can check the padding bits and discard any message not matching the expected sequence. In this case, assuming the multi-bit system outputs random codes uniformly under $\mathcal{H}_{0}$, there is still a $2^{-64+n}$ probability of ending up with the expected sequence by chance, giving $\rho\geq 2^{-64}\simeq 5e-20$. Our method benefits from comparable confidence, with $\rho=1.4e-16$ to be protected against all attacks, along with a much greater capacity.
@@ V Conclusion
In this work, we present a novel way to assert the integrity of an image by the relevant use of watermarking as a covert communication channel. Moreover we provide a definition of the image semantics, through its caption, to the recipient of the message. By using an LLM combined with an arithmetic encoder to compress the caption, the limited capacity of Hide-$\mathbb{R}$ to convey information is optimized.